OpenStudy (anonymous):

How do you solve for BOTH of the legs in a right triangle when only the hypotenuse is given?

hero (hero):

Only if both legs of a right triangle are the same length is it possible to solve it.

Parth (parthkohli):

But if you are given two angles...

Parth (parthkohli):

You can use trig if one more angle is given to ye. :-)

hero (hero):

^Then we have a different situation

hero (hero):

I guess @leximakayla should just post her problem so we can help her better.

OpenStudy (amistre64):

 if all you have is a hypot length, there are an infinite number of leg lengths that will satisfy it, but we can give a range or interval

hero (hero):

I guarantee you the user is NOT confused at this point.

OpenStudy (anonymous):

then the legs are the same length

Parth (parthkohli):

Not really. The legs are of the same length if and only if we have two 45 degree angles.

OpenStudy (anonymous):

hold on ill post a pic

hero (hero):

Yes, I was just about to suggest that...again

OpenStudy (anonymous):

number 16

hero (hero):

:)

Parth (parthkohli):

OK, that one is simply the Pythagorean Theorem.

Parth (parthkohli):

Except the 16th.

hero (hero):

No, you can still use Pythagorean Theorem for the 16th

Parth (parthkohli):

Actually, the 16th one too.

Parth (parthkohli):

Yeah. :-P

OpenStudy (anonymous):

yes, but how do you use the pythagorean theorem when both legs are missing? thats the part i dont understand

Parth (parthkohli):

Both legs are of the same length.

Parth (parthkohli):

\[x^2 + x^2 = 6^2\]

hero (hero):

\[a^2 + b^2 = c^2\] a = x = b c = 6 \[x^2 + x^2 = 6^2\] \[2x^2 = 36\] Finish solving for x

OpenStudy (anonymous):

\[x = \sqrt{18}\]

hero (hero):

Which equals \[\sqrt{9 \times 2}\] which equals \[\sqrt{9} \times \sqrt{2}\] Which equals...?

OpenStudy (anonymous):

\[3\sqrt{2?}\]

hero (hero):

Well, without the question mark under the radical :P

OpenStudy (anonymous):

i didnt mean to put that :)